OpenStudy (anonymous):

If s= 7cm and t=24cm what is the length of R?

OpenStudy (anonymous):

25 cm or 28 cm or 31cm

sam (.sam.):



sam (.sam.):

Pythagorean theorem, \[A^2=B^2+C^2\]

OpenStudy (anonymous):

is it 25cm 28cm or 31cm @.Sam.

OpenStudy (anonymous):

use \[a^2 + b^2 = c^2\] where a and b are the legs, and c is the hypotenuse.

OpenStudy (anonymous):

is it 25 cm or 28cm or 31cm? @some_someone

OpenStudy (anonymous):

are you not going to give it a try? haha well we aren't here to just give you the answer!

OpenStudy (anonymous):

ok imma try @some_someone

sam (.sam.):

You have \[A^2=B^2+C^2\] Substitute in you got \[A=\sqrt{7^2+24^2}\] Solve

OpenStudy (anonymous):

is it 31cm? @some_someone @.Sam.

sam (.sam.):

Nope

OpenStudy (anonymous):

28 then

OpenStudy (anonymous):

28 @.Sam.

sam (.sam.):

\[A=\sqrt{7^2+24^2}\] What do you get from here?

OpenStudy (anonymous):

28 @.Sam.

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] \[7^2 + 24^2 = c^2\] \[49 + 576 = c^2\] \[625 = c^2\] \[\sqrt{625} = \sqrt{c^2}\] \[25 = c\]